# Not Everyone Wins with LLMs: Behavioral Patterns and Pedagogical Implications in AI-assisted Data Analysis

### v1

@@ 7.2. What should we teach on AI use skills
Table 6. Success and failure AI use behaviors with corresponding missing AI use skills.
@@
Dispositional.
The lack of persistence may also result in early stops. For instance, when a student tried to add labels on top of a pie chart, the generated code produced errors, and the fixed code, when simply pressing the “explain error” button, did not fix the issue. The student then stopped trying, accepting the basic pie chart without the label and commented “seems like it either do the label or a pie” (HW2 think-aloud). In comparison, a success case will display resilience, where the student keeps trying to refine outputs until they achieve the intended goals. For example, as shown in a sequence of prompt iterations, the student persisted through different AI mistakes to produce the ideal visualization for EDA:
$\rightarrow$ regenerate to make the grouped bar chart
$\rightarrow$ Change this to output one such chart per year. There should be one chart for every year in the range
(HW2 log).
8. Discussion and Future Work
8.1. Design and Pedagogy Implications
Make AI mental models explicit.
Unlike simple chat interfaces, AI integrated in a notebook exposes multiple affordances (code cells, outputs, error panes, menu buttons, etc.). We observed in the study that students transferred assumptions from more capable general-purpose AIs to their Colab working environment, miscalibrating what the embedded model could access or do. This produced confusion about AI task space (what to ask) and context space (what the model can “see”). As shown in Table 5, a novice during think-aloud prompted the model “Don’t use nonexistent columns, use what you have in the datasheet” after three repeated KeyError and commented “Perplexity can do that,” assuming that Colab Gemini would have access to the imported dataframe like other systems. It took them more than five repeated failures to realize that they should adjust their mental model for the system at hand, highlighting the need for explicit conceptual models and metacognitive updating abilities. On the design side, systems with GenAI assistants should make their task and context scope more legible (e.g., what files and variables are visible to GenAI, what operations are supported, etc). Teaching “what to ask this specific AI” and “what this specific AI knows” is likely getting obsolete as models and systems rapidly evolve. More importantly, instructions are needed for metacognitive abilities to quickly adjust mental models to different AI system capabilities.
As we found in Section 7.1, using AI to gain additional, evaluative information was a consistent challenge. Asking AI to explain code, bugs, or critique outputs remained limited, despite its potential benefit to improve understanding. We especially observe that students may not persist in reading convoluted output, reflecting reluctance or difficulty in engaging with evaluative or complex information. While there are one-click affordances (explain error buttons) that aim to make it easier to start evaluation, the cognitive load of seeing a lengthy error explanation often creates resistance or reduces the chance for students to learn critical evaluation skills.
8.2. Limitations and Future Work
Our analyses suggest that students’ technical expertise significantly predicts performance during LLM use (similar to prior literature (Chen et al., 2024)), whereas LLM and communication expertise factors did not (Section 5). Other work also shows that LLMs can narrow gaps between novices and experts (Pickering et al., 2025), or even impede expert performance under certain conditions (Becker et al., 2025). This highlights that technical expertise is not the only factor shaping effective LLM use. Our log analyses revealed the importance of AI use strategies, which point toward the need to define and evaluate “AI use expertise” more directly, as what we have attempted in Section 7.2.
To probe the interplay between technical expertise and LLM use, we compared performance between experts and novices in the timed HW0 under conditions with and without LLMs (Table 1). A Mann–Whitney U test (as data violated normality assumptions) (McKnight and Najab, 2010) showed a significant expertise gap when students did not use LLMs ($p=0.027<0.05$), but the gap disappeared when LLM was used ($p>0.05$). This suggests that under time pressure, LLMs may temporarily offset expertise advantages, indicating that factors such as time constraints also influence the effectiveness of LLM use. We also measured affective factors (motivation, self-efficacy) in self-reported survey items (Appendix B), but none significantly predicted performance.
Our sample included 36 students in authentic classroom contexts. While this ensured ecological validity, it is a small sample and constrained our ability to isolate factors such as undisclosed external tool use (e.g., ChatGPT alongside Colab Gemini) or overlapping expertise dimensions. For example, Python and data science proficiency were moderately correlated in our sample, so we could not examine their independent effects on effective LLM use. Future studies should recruit distinct participant groups or employ pretests. More robust performance-based assessment and controlled experiments are needed to disentangle the interplay between more nuanced factors such as domain expertise, implementation expertise, AI-use expertise, affect, and situational factors like time pressure.
Going broader: extend to more diverse task contexts.
While the learning objectives for AI use skills (Table 4) were derived from programmatic data analysis tasks, they should extend beyond our study context. Many of the objectives are broadly relevant to problem-solving and higher-order thinking skills; for instance, clarifying the task space of any tool, communicating requirements effectively with decomposition, and verifying results. Future research could examine whether instruction around AI use transfers to these more general competencies, and explore how to evaluate AI use expertise more systematically (as self-rated LLM expertise may not be appropriate).
Finally, we chose Google Colab for the class because it embeds a free, in-house Gemini with different UIs, which enabled us to explore diverse behavioral AI use patterns in our research questions. However, our data collection relied on a custom logging script in the browser, which can be less robust than existing Jupyter notebook logging toolkits like JELAI (Valle Torre et al., 2025). For example, our approach may capture an auto-completion as users’ code edits (although infrequent), and Colab’s frequently changing interface raises concerns about long-term replicability. Future work could build on established logging toolkits to improve data quality and generalizability.
9. Conclusion
LLMs promise to broaden access to technical work, yet our study shows that not everyone benefits equally. Technical expertise remained the strongest predictor of success in homework, and experts engaged with AI more strategically than novices. Behavioral log analysis highlighted that experts provide clearer context, decompose tasks, and strategically use AI, while novices tend to rely on AI reactively in the face of challenges. While demonstrations and extended time helped students adopt clearer prompting practices, evaluative skills remained challenging. These findings suggest that effective AI use is not just about producing good prompts, but also more about cultivating durable competencies in multiple dimensions. We translate these insights into design and pedagogy implications for AI literacy: explicate mental models, clarify before generating, and scaffold verification. Ultimately, effective AI use is not just better prompts — it is about better problem solving.
Acknowledgements.
Thanks to all the participants for the study, and thanks to all LearnLab members who provided feedback on this work. Thanks to the National Science Foundation (award CNS-2213791, 2414915) and Google Academic Research Award for partial support of this work. Acknowledgment of LLM Use. We acknowledge the use of ChatGPT to assist with language editing, summarization, and restructuring of text during manuscript preparation. Claude was used to assist the analysis of log data; implementation details are included in Section 6.1.2, prompts and settings are attached in Appendix D. Copilot was used to assist the visualization and code writing to process and clean our dataset, and Gemini in Google Colab was used to prepare the homework materials and reference solutions.
@@ Appendix A Codebook for task challenge and AI use success and failure behaviors
Form
Input
Appendix B Survey Instruments
@@
B.1. Pre Survey Only Questions
@@
B.2. Shared Questions Across Post Surveys
@@
B.3. HW0 Post Survey Only Questions
@@
B.4. HW1–HW4 Post Survey Only Questions
@@
B.5. Final Post Survey Only Questions
@@
Appendix C Composite Competency Inclusions
Technical competency (composite of programming, data science, and tools)
@@ Appendix C Composite Competency Inclusions
Appendix D Prompt for Log Analysis
@@ D.1. Segmentation Prompt
Decision boundary for different intent (and split into new episode): a different EDA question is a different intent a different plot is a different intent, even if it visualizes the same relationship or answers the same EDA question cleaning a different column is a different intent removing outliers is different intent from making a column dtype numeric, even if they are both cleaning the same column Episode Steps: each episode includes the following four steps; early stops/failures may not reach beyond step 1: (1) Form Intent: Deciding what needs to be done to progress the task (e.g., remove outliers in a column, create a bar chart, refine a visual by sorting by count, clarify concept, break down a complex goal). Each unique intent should start a new episode. Writing comments and prompting model are common behaviors that can be used to infer intent. (2) Form Input: Translating intent into input (Python code in edit_code or an AI prompt in converse or generateCode) that can generate the desired result. (3) Understand Output: Reading and interpreting the system’s response (llm output in event_response, executed_output, error messages in executed_error, plots, tables, or text). (4) Assess Output: Evaluating whether the output is correct, useful, and aligned with the intended goal. Instructions:
@@ 1. Segment logs into episodes:
Start a new episode when you observe a clear change in intent (see decision boundary above), OR if there is a time gap $\geq$ 15 minutes between consecutive events (use datetime_timestamp). Ensure all log events are covered. Episodes should be consecutive, from the first log to the last. Prefer fewer coherent episodes unless there is a clear goal shift or long gap. Each episode must have timestamp_start and timestamp_end taken from the episode’s first/last event.
2. Identify the episode’s overall intent: Infer from code, prompts, and outputs (e.g., remove outliers in a column, create a bar chart, refine a visual by sorting by count). Write 1–2 sentences describing what the user was trying to do and what happened, inferred from the log events. Output Format (JSON array of episode objects):
@@
D.2. Step-wise Annotation Prompt
@@
(1) Intent – forming a correct goal (2) Input – expressing the goal clearly in code or prompt to generate desired output (3) Understand – interpreting output/errors (4) Assess – judging output correctness & alignment with goal
Users may display different AI-use behaviors in different steps of the episode: intent-step: asks AI to explain concepts or suggest steps; asks AI to break down steps. ai_explain_concepts ai_suggest_steps_or_plan ai_breakdown_intent input-step: asks AI to improve/clarify the prompt; ask AI to edit code or generate code ai_improve_prompt ai_edit_partial_code ai_generate_code understand-step: asks AI to explain code or error, press_explain_error, press_explain_code, or clarifies APIs ai_explain_bug_or_error ai_explain_code_or_api assess-step: asks AI to critique output; queries alignment with goals/insight quality ai_critique_output step-agnostic-ai-use: good AI use can happen across steps (whenever the user is using AI) clear_instruction_in_prompt provide_context_in_prompt decompose_task_in_prompt Behavior Examples
@@
Some example behaviors that demonstrated challenges in each step: intent-step: user stuck (“don’t know what to do”), failure to form intent (e.g., “what else can I do?” / “give me steps”), wrong intent (e.g., asks for adding “\$” instead of making a price column numeric), cannot decompose a complex goal. input-step: vague pronouns (“fix it”) without shared context, missing/incorrect context leading to hallucinations (e.g., KeyError of made-up columns), faulty code (syntax error, wrong API usage), edit_code followed by executed_error. understand-step: copy-paste generated code blindly, immediate “fix it” after press_explain_error without reading, repeating error loops with no adjustment. assess-step: accepts generation without verification (very short intervals), keeps ineffective chart/insight misaligned with goals. Some success behaviors for each step: intent-step: final intent is clear and correct (e.g., valid cleaning procedure, valid EDA question) input-step: input code/prompt are clear and correct (errors resolved) understand-step: demonstrated understanding via a concrete next action or time spent reading assess-step: result (data/chart/stats) is correct; or student requests targeted improvement based on assessment Instructions
@@ 1. For each step of the episode, trace key events and summarize into Actions & Returns with evidence:
Action: initiate an attempt (e.g., edit code, enter prompt, send Gemini prompt, UI button like execute after editing code, press_explain_error, etc.) Return: immediate observable system feedback (stdout/stderr, exceptions, plots, metrics, AI response) Steps and evidence event_ids must be in chronological order.
2. For each step, annotate if a challenge was encountered, and whether AI was used strategically to overcome it: Use the challenge/success behavior examples to mark encountered_challenge. If AI was used well, tag behaviors (e.g., ai_explain_concepts, clear_instruction_in_prompt) and briefly explain how it helped. If not overcome, note which AI-use behaviors could have helped and why. 3. Decide if the step was successful, and explain: step_success can be true/false (partial progress or step not reached counts as false). Provide 2–3 sentences citing how the outcome is inferred from logs; include representative event_names and short excerpts from event_content/errors/prompts or relevant timestamp gaps. Steps can be independent (e.g., wrong intent but correct understanding/assessment). Visual-output assessment
@@
(1) Identify if there is visualization code (e.g., alt.Chart) intended to answer an EDA question. Evaluate the code even if the actual image is not available. (2) Decide if the visualization is appropriate. If not, did the user refine it appropriately? Examples of assessment failures include: accepting an overly-busy chart (e.g., all genres when only top genres matter); using the wrong plot type (e.g., line instead of scatter for correlation); using a boxplot to show average rating (producing short lines rather than showing a range).

### v2

@@ Appendix A Codebook for task challenge and AI use success and failure behaviors
Table 7. Mapping between the success and failure AI use behaviors (code frequency in Figure 9) and corresponding missing AI use skills (learning objectives).
@@
Examples linking log excerpts to specific AI-use learning objectives, illustrating conceptual, procedural, metacognitive, and dispositional challenges.
@@
Appendix B Study Method Details
Here we list the data source details for each analysis method (Section 4):
  • RQ1 (Section 5):
    • Mann-Whitney U test (36 students, HW0, 36 submissions),
    • LMER (36 students, HW1-4, 144 submissions),
  • RQ2 (Section 6):
    • AI use behaviors codebook (Table 2) construction (10 students, 5 screen videos + logs from HW0, 5 think-aloud videos + logs from HW1-4),
    • Annotated log analysis (28 students, HW0-2, 36 submissions),
    • See method details and justifications in Section 4.3,
  • RQ3 (Section 7):
    • AI use skills codebook (Table 4) construction (36 students, HW0-4, 180 post-surveys),
    • Annotated log analysis (28 students, HW0-2, 36 submissions),
  • Qualitative examples could be selected from all students logs / videos / surveys from all homework.
Figure 9. Distribution of AI use code frequency averaged per student, aggregated across all steps and clustered by AI usage code types. Refer to Table 2 for definitions.
Bar charts showing the overall frequency of each AI usage behavior across logs. Providing context and requesting steps are frequent, while prompt refinement and partial code edits are rare.
Appendix C Survey Instruments
@@
C.1. Pre Survey Only Questions
@@
C.2. Shared Questions Across Post Surveys
@@
C.3. HW0 Post Survey Only Questions
@@
C.4. HW1–HW4 Post Survey Only Questions
@@
C.5. Final Post Survey Only Questions
@@
Appendix D Composite Experience Measures
D.1. Grouping Construction Process
The experience variables are rationally constructed to enable interpretability and are also empirically justified with correlation analysis and principal component analysis (PCA) (Greenacre et al., 2022). As shown in Figure 10a, PC1 captures the most variance ($\sim$49%), aligning with technical experience; PC2 explains additional variance related to LLM experiences ($\sim$11%); Communication-related measures cluster around the negative region of both PC1 and PC2. Note that given the low sample size and limited statistical power, we used PCA only to guide and visually verify the clustering of variables. We further validated the independence of these experience variables using correlation analysis and varience inflation factors (VIF) (Marcoulides and Raykov, 2019), as shown in Figure 10b. The lack of multicollinearity enabled further regression analysis of the experience variables in Section 5.
(a) PCA of self-reported experience measures.
PCA scatterplot of self-reported experience measures, with each variable positioned by its loading on PC1 and PC2. Technical experience variables cluster along the positive PC1 axis, LLM-related variables along positive PC2, and communication measures in the negative regions of both axes.
(b) Pairwise-correlation matrix and VIF scores for composite experience measures.
Tech LLM Commun. VIF
Tech 1.00 0.41 -0.23 1.22
LLM 0.41 1.00 -0.28 1.26
Commun. -0.23 -0.28 1.00 1.10
Variance Inflation Factor analysis and pairwise correlation matrix showing low multicollinearity among technical, LLM, and communication experience variables.
Figure 10. Overview of self-reported experience measures: (a) PCA and (b) pairwise correlations and VIF scores.
D.2. Grouping Items
Technical experience (composite of programming, data science, and tools)
@@ D.2. Grouping Items
Appendix E Prompt for Log Analysis
@@ E.1. Segmentation Prompt
Decision boundary for different intent (and split into new episode):
  • a different EDA question is a different intent
    • a different plot is a different intent, even if it visualizes the same relationship or answers the same EDA question
  • cleaning a different column is a different intent
    • removing outliers is different intent from making a column dtype numeric, even if they are both cleaning the same column
Episode Steps: each episode includes the following four steps; early stops/failures may not reach beyond step 1:
  1. (1)
    Form Intent: Deciding what needs to be done to progress the task (e.g., remove outliers in a column, create a bar chart, refine a visual by sorting by count, clarify concept, break down a complex goal). Each unique intent should start a new episode. Writing comments and prompting model are common behaviors that can be used to infer intent.
  2. (2)
    Form Input: Translating intent into input (Python code in edit_code or an AI prompt in converse or generateCode) that can generate the desired result.
  3. (3)
    Understand Output: Reading and interpreting the system’s response (llm output in event_response, executed_output, error messages in executed_error, plots, tables, or text).
  4. (4)
    Assess Output: Evaluating whether the output is correct, useful, and aligned with the intended goal.
Instructions:
@@ 1. Segment logs into episodes:
  • Start a new episode when you observe a clear change in intent (see decision boundary above), OR if there is a time gap $\geq$ 15 minutes between consecutive events (use datetime_timestamp).
  • Ensure all log events are covered. Episodes should be consecutive, from the first log to the last.
  • Prefer fewer coherent episodes unless there is a clear goal shift or long gap.
  • Each episode must have timestamp_start and timestamp_end taken from the episode’s first/last event.
2. Identify the episode’s overall intent:
  • Infer from code, prompts, and outputs (e.g., remove outliers in a column, create a bar chart, refine a visual by sorting by count).
  • Write 1–2 sentences describing what the user was trying to do and what happened, inferred from the log events.
Output Format (JSON array of episode objects):
@@ 2. Identify the episode’s overall intent:
E.2. Step-wise Annotation Prompt
@@
  1. (1)
    Intent – forming a correct goal
  2. (2)
    Input – expressing the goal clearly in code or prompt to generate desired output
  3. (3)
    Understand – interpreting output/errors
  4. (4)
    Assess – judging output correctness & alignment with goal
Users may display different AI-use behaviors in different steps of the episode:
  • intent-step: asks AI to explain concepts or suggest steps; asks AI to break down steps.
    • ai_explain_concepts
    • ai_suggest_steps_or_plan
    • ai_breakdown_intent
  • input-step: asks AI to improve/clarify the prompt; ask AI to edit code or generate code
    • ai_improve_prompt
    • ai_edit_partial_code
    • ai_generate_code
  • understand-step: asks AI to explain code or error, press_explain_error, press_explain_code, or clarifies APIs
    • ai_explain_bug_or_error
    • ai_explain_code_or_api
  • assess-step: asks AI to critique output; queries alignment with goals/insight quality
    • ai_critique_output
  • step-agnostic-ai-use: good AI use can happen across steps (whenever the user is using AI)
    • clear_instruction_in_prompt
    • provide_context_in_prompt
    • decompose_task_in_prompt
Behavior Examples
@@
Some example behaviors that demonstrated challenges in each step:
  • intent-step: user stuck (“don’t know what to do”), failure to form intent (e.g., “what else can I do?” / “give me steps”), wrong intent (e.g., asks for adding “\$” instead of making a price column numeric), cannot decompose a complex goal.
  • input-step: vague pronouns (“fix it”) without shared context, missing/incorrect context leading to hallucinations (e.g., KeyError of made-up columns), faulty code (syntax error, wrong API usage), edit_code followed by executed_error.
  • understand-step: copy-paste generated code blindly, immediate “fix it” after press_explain_error without reading, repeating error loops with no adjustment.
  • assess-step: accepts generation without verification (very short intervals), keeps ineffective chart/insight misaligned with goals.
Some success behaviors for each step:
  • intent-step: final intent is clear and correct (e.g., valid cleaning procedure, valid EDA question)
  • input-step: input code/prompt are clear and correct (errors resolved)
  • understand-step: demonstrated understanding via a concrete next action or time spent reading
  • assess-step: result (data/chart/stats) is correct; or student requests targeted improvement based on assessment
Instructions
@@ 1. For each step of the episode, trace key events and summarize into Actions & Returns with evidence:
  • Action: initiate an attempt (e.g., edit code, enter prompt, send Gemini prompt, UI button like execute after editing code, press_explain_error, etc.)
  • Return: immediate observable system feedback (stdout/stderr, exceptions, plots, metrics, AI response)
  • Steps and evidence event_ids must be in chronological order.
2. For each step, annotate if a challenge was encountered, and whether AI was used strategically to overcome it:
  • Use the challenge/success behavior examples to mark encountered_challenge.
  • If AI was used well, tag behaviors (e.g., ai_explain_concepts, clear_instruction_in_prompt) and briefly explain how it helped.
  • If not overcome, note which AI-use behaviors could have helped and why.
3. Decide if the step was successful, and explain:
  • step_success can be true/false (partial progress or step not reached counts as false).
  • Provide 2–3 sentences citing how the outcome is inferred from logs; include representative event_names and short excerpts from event_content/errors/prompts or relevant timestamp gaps.
  • Steps can be independent (e.g., wrong intent but correct understanding/assessment).
Visual-output assessment
@@ 3. Decide if the step was successful, and explain:
  1. (1)
    Identify if there is visualization code (e.g., alt.Chart) intended to answer an EDA question. Evaluate the code even if the actual image is not available.
  2. (2)
    Decide if the visualization is appropriate. If not, did the user refine it appropriately? Examples of assessment failures include:
    • accepting an overly-busy chart (e.g., all genres when only top genres matter);
    • using the wrong plot type (e.g., line instead of scatter for correlation);
    • using a boxplot to show average rating (producing short lines rather than showing a range).